# A New Perspective on Double-S-Curve Motions of Higher Order and Optimal Motion Planning

### v1

@@ 5.2 Jerk-Controlled Motions
$\hat{T}_{0}$ $=2\hat{T}_{1}$ $(\hat{a}_{\min}\mathrm{curve})$ (66)
$\hat{T}_{1}$ $=2\hat{T}_{2}$ $(\hat{v}_{\min}\mathrm{curve})$ (67)
Figure 4: Contour plot of $\hat{T}$ as a function of $\hat{a}$ and $\hat{v}$ according to Equation (63). The global minimum (blue cross) is located at $\hat{a}_{\star}=\sqrt[3]{\frac{1}{2}}$, $\hat{v}_{\star}=\sqrt[3]{\frac{1}{4}}$ and $\hat{T}_{\star}=\sqrt[3]{32}$. The red solid boundary curves of the admissible region are given by $\hat{v}=\hat{a}^{2}$ ($\hat{v}_{\min}$ curve) and $\hat{v}=-\frac{1}{2}+\sqrt{\frac{1}{4}}\hat{a}^{4}+\hat{a}$ ($\hat{a}_{\min}$ curve).
If such a condition is satisfied the associated order has no cruising phase. For example, $\hat{T}_{1}=2\hat{T}_{2}$ means that $\ddot{x}$ has a triangular shape if it is not equal to 0 (see Figure 2 both upper panels). If there is no cruising phase in $\ddot{x}$, the time for reading $\hat{v}$ (equal to $\hat{T}_{1}$) is exactly twice the time needed for reading $\hat{a}$ (equal to $\hat{T}_{2}$). This corresponds to the fact that for every acceleration a certain velocity is quasi automatically achieved. Therefore, not all of the ($\hat{v},\hat{a}$)-paris are actually possible. And it is precisely these borderline cases that are reflected in the conditions (66,67).
Furthermore, the global minimum is determined by simultaneously satisfying the equations (66,67). It is located at $\hat{a}_{\star}=\sqrt[3]{\frac{1}{2}}$, $\hat{v}_{\star}=\sqrt[3]{\frac{1}{4}}$ and $\hat{T}_{\star}=\sqrt[3]{32}$.
Let us now take a closer look at MTT. The four possible cases discussed in Section 3.2 can be easily recoveres and re-interpreted within the fish. As above, $\hat{v}_{\max}$ and $\hat{a}_{\max}$ are horizontal and vertical cut-offs, respectively. Then:
  • 4-phase case: $\hat{v}_{\max}\geq\hat{v}_{\star}$ and $\hat{a}_{\max}\geq\hat{a}_{\star}$ (no active bound), and the solution of MTT is euqivalent to the global minimum (see Figure 5, upper left panel).
  • 5-phase case: $\hat{v}_{\max}<\hat{v}_{\star}$ and $\hat{a}_{\max}\geq\hat{a}_{\star}$ (velocity bound active), and the solution of MTT is given by te intersection of $\hat{v}_{\max}$ and the $\hat{a}_{\min}$ curve [^5] Note that the inactive acceleration bound is equivalent to Equation (66). see Figure 5, upper right panel.
  • 6-phase case: $\hat{v}_{\max}\geq\hat{v}_{\star}$ and $\hat{a}_{\max}<\hat{v}_{\star}$ (acceleration bound active), and the solution of MTT is given by the intersection of $\hat{a}_{\max}$ and the $\hat{v}_{\min}$ curve [^6] The inactive velocity bound is equivalent to Equation (67)., see Figure 5, lower left panel.
  • 7-phase case: $\hat{v}_{\max}<\hat{v}_{\star}$ and $\hat{a}_{\max}<\hat{a}_{\star}$ (both bounds active); then $\hat{v}=\hat{v}_{\max}$ and $\hat{a}=\hat{a}_{\max}$ and Equation (59) becomes
@@
      Otherwise, the solution is determined by the intersection of the fixed contour curv and the vertical cut-off defined by $\hat{a}_{\max}$ (see Figure 6).
@@
All in all, the relevant optimal values can be obtained by targeted solving of algebraic equations. In addition, the fish diagram provides an overview which case actually occurs, i.e. of how many phases the motion consists and which particular bounds are active. Noteworthy, independent of the case MTT or MDT, Equation (59) remains overall valid.
@@
Figure 5: Schematic overview of the 4 possible cases of a jerk-controlled motion distinguished by the number of phases. The solution of MTT is marked with a green bullet (except for the 4-phase case, where the solution is equivalent to the global minimum highlighted by the blue cross). In each case, the admissible region lies below and left the $\hat{v}_{\max}$ and $\hat{a}_{\max}$ bounds, respectively, and is additionally bounded ba the red curves.
Figure 6: F4, also Figure 6 kommt noch
Figure 7: F5, also Figure 7 kommt noch
@@ 5.3 Snap-Controlled Motions
$\hat{T}$ $=Tp_{\max}^{\frac{1}{4}}s^{-\frac{1}{4}}\quad\hat{v}$ $=vp_{\max}^{\frac{1}{4}}s^{-\frac{1}{4}}$ (73)
$\hat{a}$ $=ap_{\max}^{\frac{1}{2}}s^{-\frac{3}{2}}\quad\hat{\jmath}$ $=jp_{\max}^{\frac{3}{4}}s^{-\frac{7}{4}}$
@@
allows us to plot a three-dimensional fish (see Figure 8). The contour surfaces of constant $\hat{T}$ values are indicated by the colored curves circulating around the fish. The conditions
@@
$\hat{\jmath}_{\star}=\sqrt[4]{\frac{1}{8}},\quad\hat{a}_{\star}=\sqrt{\frac{1}{8}},\quad\hat{v}_{\star}=\sqrt[4]{\frac{1}{32}},\quad\hat{T}_{\star}=\sqrt[4]{512}.$ (75)
@@
  • the bounds $\hat{v}_{\max}$, $\hat{a}_{\max}$ and $\hat{j}_{\max}$ are now cutting planes.
  • If no bound is active (i.e. the solution of MTT s equivalent to the global minimum), $\dddot{x}(t)$ is “fully triangular”, and none of the orders have a cruising phase. Hence this kind of motion consists of 8 phases which is the minimum number of phases for a snap-controlled motion.
@@
  • This distinction can be made systematically by switching the conditions of (74) on or off independently from each other and counting how independently from each other and counting how many combinations result (indeed, $2^{3}=8$).
@@
Figure 8: Three-dimensional plot of $\hat{T}$ as a function of $\hat{v}$, $\hat{a}$ and $\hat{\jmath}$ according to Equations (69,73). The surfaced of constant $\hat{T}$ value are indicated by the colored curves (from orange for high $\hat{T}$ to dark purple for low $\hat{T}$), where the actual surface embraced by the curves is not shown for a better visibility of the entire figure. The three red curves frame the region of all admissible combinations of ($\hat{v},\hat{a},\hat{\jmath}$). hey are obtained by combining two of the three conditions of (74), e.g., $\hat{T}_{1}=2\hat{T}_{2}$ and $\hat{T}_{2}=2\hat{T}_{3}$ fix the red olid curve. A single condition like $\hat{T}_{1}=2\hat{T}_{2}$ corresponds to a surrounding surface (not shown). The global minimum which is determined by the simultaneous consideration of all conditions of (74) is marked by the blue cross.
@@ 5.4 Systematics for Arbitrary Order
As already mentioned, such a motion has at least 8 phases and at most 15 phases. The click already mentioned at the end of Section 5.3 is the 5th order $x^{(5)}(t)$. Setting $c=x_{5}=\max\limits_{t\in[0,T]}|x^{(5)}(t)|$ a achieved click leads to
@@
Also even higher order, for example the yeet $x^{(6)}(t)$ and the pow $x^{(7)}(t)$, can be considered. We will continue with an arbitrary order $N$. The dimensional transformation to obtain $\hat{x}_{0}=\hat{x}_{N}=1$ is in general given by
$\hat{x}_{n}$ $=x_{n}x_{0}^{-(1+\frac{n}{N})}x_{N}^{\frac{n}{N}}$ (81)
$\hat{T}$ $=Tx_{0}^{-\frac{1}{N}}x_{N}^{\frac{1}{N}}$ (82)
@@
$\hat{x}_{N-1}$ $=2^{\frac{1}{2N}(N-1)(2-N)}$ (83)
$\hat{x}_{n}$ $=\hat{x}_{N-1}^{N-n}2^{\frac{1}{2}(n-N-1)(2-N)}$ (84)
@@
Clearly, $\hat{T}_{N-1}=\hat{x}_{N-1}$. Due to Equation (77) it is easy to find that
$\hat{T}_{n}=2^{N-n-1}\hat{x}_{N-1}$ (87)
@@
as a recursion for $h$ (with h(N1)=0h(N-1)=0). Then the solution is $h(n)=\frac{1}{2}(n-N+1)(n-N+2)$ which proves Equation (84). The condition $\hat{T}_{0}=2\hat{T}_{1}$ which is equivalent to $\frac{1}{\hat{x}_{1}}=\hat{T}_{1}$ finally determines $\hat{x}_{N-1}$ by using Equations (84) and (86). Equation (86) is a direct implication. ∎
The following table 1 lists some examples.
N $\hat{T}_{0}$ $\hat{v}$ $\hat{a}$ $\hat{\jmath}$
@@ 4 $\sqrt[4]{512}$ $\sqrt[4]{\frac{1}{32}}$ $\sqrt{\frac{1}{8}}$ $\sqrt[4]{\frac{1}{8}}$
it can seen from Equation (85) that the minimal time grows exponentially with the order of the motion [^7] While for the jerk-controlled motion $\hat{T}_{0}\approx 3.17$, the case $N=6$ (yeet-controlled motion) and $N=7$ (pow-controlled motion) yield $\hat{T}_{0}\approx 10.08$ and $\hat{T}_{0}\approx 14.49$.. So the time intervals, where the highest order is non-vanishing, become rapidly shorter and the ratios $\frac{x_{n}}{x_{n+1}}$ for higher $n$ shrink.
All in all, the impact of higher order like snap, click, yeet and pow is rather small (especially for the most practical purposes).
@@ 6 Summary and Outlook
for jerk-controlled motions, is valid for any number of phases the motion can have. Furthermore, arbitrarily high motion orders can be considered. This allows for a reinterpretation and solution of well-known standard trajectory optimization problems (such as motions with minimum-time or motions with minimum velocity). In particular, the condition $T_{n}=2T_{n+1}$ enables the determination of the active constraints for th respective case. It can also be estimated what effect considering higher orders has on the overall time horizon. Therefore, it seems surprising that – although the equation is not exceptionally complex and covers a large number of cases – it hat not previously been presented in this form in the literature.
Besides the practical case of application, it should be noted that motions with the minimum velocity are also energy-optimal in many cases, because a long phase without acceleration combined with a short phase of higher acceleration outweights motions motions with long phases of lower acceleration.
Clearly, many extensions exist. First, the zero boundary condition can be relaxed. Furthermore, it would be possible to consider other classes of trajectories (XXX). The expected results would then likely be considerably more unwiddly than those obtained here.
Another possibility – particularly mathematically interesting – lies in the limit $N\rightarrow\infty$ of infinite order. This raises questions of convergence (see the series in Equation (32) in Theorem 5) as well as questions concernung the relationship with $\mathcal{C}^{\infty}$ functions and mollifiers. Due to the infinitely large number of intervals in a finite time interval, it is also worthwhile to consider methods of fractal geometry (such as fractal strings, see [Lapidus]). However, all these investigations of interesting connections are left for future work.

### v2

@@ 5.2 Jerk-Controlled Motions
$\hat{T}_{0}$ $=2\hat{T}_{1}\quad(\hat{a}_{\min}\penalty 10000\ \mathrm{curve})$ (58)
$\hat{T}_{1}$ $=2\hat{T}_{2}\quad(\hat{v}_{\min}\penalty 10000\ \mathrm{curve})$ (59)
Figure 4: Contour plot of $\hat{T}$ as a function of $\hat{a}$ and $\hat{v}$ according to Equation (55). The global minimum (blue cross) is located at $\hat{a}_{\star}=\sqrt[3]{\frac{1}{2}}$, $\hat{v}^{\star}=\sqrt[3]{\frac{1}{4}}$ and $\hat{T}^{\star}=\sqrt[3]{32}$. The red solid boundary curves of the admissible region are given by $\hat{v}=\hat{a}^{2}$ ($\hat{v}_{\min}$ curve) and $\hat{v}=-\frac{1}{2}\hat{a}^{2}+\sqrt{\frac{1}{4}\hat{a}^{4}+\hat{a}}$ ($\hat{a}_{\min}$ curve).
If such a condition is satisfied the associated order has no cruising phase. For example, $\hat{T}_{1}=2\hat{T}_{2}$ means that $\ddot{x}$ has a triangular shape if it is not equal to 0 (see Figure 2 both upper panels). If there is no cruising phase in $\ddot{x}$, the time for reaching $\hat{v}$ (equal to $\hat{T}_{1}$) is exactly twice the time needed for reaching $\hat{a}$ (equal to $\hat{T}_{2}$). This corresponds to the fact that for every acceleration a certain velocity is quasi automatically achieved. Therefore, not all of the ($\hat{v},\hat{a}$)-pairs are actually possible. And it is precisely these borderline cases that are reflected in the conditions (58, 59).
Furthermore, the global minimum is determined by simultaneously satisfying the equations (58, 59). It is located at $\hat{a}^{\star}=\sqrt[3]{\frac{1}{2}}$, $\hat{v}^{\star}=\sqrt[3]{\frac{1}{4}}$ and $\hat{T}^{\star}=\sqrt[3]{32}$.
Let us now take a closer look at MTT. The four possible cases discussed in Section 3.2 can be easily recovered and reinterpreted within the fish diagram. As above, $\hat{v}_{\max}$ and $\hat{a}_{\max}$ are horizontal and vertical cut-offs, respectively. Then:
  • 4-phase case: $\hat{v}_{\max}\geq\hat{v}^{\star}$ and $\hat{a}_{\max}\geq\hat{a}^{\star}$ (no active bound), and the solution of MTT is euqivalent to the global minimum (see Figure 5, upper left panel).
  • 5-phase case: $\hat{v}_{\max}<\hat{v}^{\star}$ and $\hat{a}_{\max}\geq\hat{a}^{\star}$ (velocity bound active), and the solution of MTT is given by the intersection of $\hat{v}_{\max}$ and the $\hat{a}_{\min}$ curve [^7] Note that the inactive acceleration bound is equivalent to Equation (58). (see Figure 5, upper right panel).
  • 6-phase case: $\hat{v}_{\max}\geq\hat{v}^{\star}$ and $\hat{a}_{\max}<\hat{v}^{\star}$ (acceleration bound active), and the solution of MTT is given by the intersection of $\hat{a}_{\max}$ and the $\hat{v}_{\min}$ curve [^8] The inactive velocity bound is equivalent to Equation (59). (see Figure 5, lower left panel).
  • 7-phase case: $\hat{v}_{\max}<\hat{v}^{\star}$ and $\hat{a}_{\max}<\hat{a}^{\star}$ (both bounds active); then $\hat{v}=\hat{v}_{\max}$ and $\hat{a}=\hat{a}_{\max}$ (see Figure 5, lower right panel) and Equation (51) becomes
@@
      Otherwise, the solution is determined by the intersection of the fixed contour curve and the vertical cut-off defined by $\hat{a}_{\max}$ (see Figure 6, right panel).
@@
All in all, the relevant optimal values can be obtained by targeted solving of algebraic equations. In addition, the fish diagram provides an overview of which case actually occurs, i.e. of how many phases the motion consists and which particular bounds are active. Noteworthy, independent of the cases and MTT or MDT, Equation (51) remains overall valid.
@@
Figure 5: Schematic overview of the 4 possible cases of a jerk-controlled motion distinguished by the number of phases. The solution of MTT is marked with a green bullet (except for the 4-phase case, where the solution is equivalent to the global minimum highlighted by the blue cross). In each case, the admissible region lies below and left the $\hat{v}_{\max}$ and $\hat{a}_{\max}$ bounds, respectively, and is additionally bounded by the red curves.
Figure 6: Schematic illustration of the solution of MDT with $M=1$, i.e. minimum-velocity motions. In any case, the solution is indicated by an orange bullet. Left: if $\hat{a}_{\max}>\hat{a}^{\star}$, the minimal velocity is achieved at the point where the contour curve belonging to the given time horizon $\hat{T}_{\mathrm{MDT}}$ leaves the admissible region and intersects the $\hat{v}_{\max}$ curve. Right: For small enough values of $\hat{a}_{\max}$ the contour curve reaches the acceleration boundary before the $\hat{v}_{\min}$ are. The solution is determined by this intersection.
Figure 7: Analogously depiction of MDT with $M=2$ (minimum-acceleration motions) as in Figure 6. The left panel shows the case where the solution is determined by the $\hat{a}_{\min}$ curve, whereas the right panel displays the case where the solution is determined by the $\hat{v}_{\max}$ bound.
@@ 5.3 Snap-Controlled Motions
$\hat{T}$ $=Tp_{\max}^{\frac{1}{4}}s^{-\frac{1}{4}}$ (65)
$\hat{v}$ $=vp_{\max}^{-\frac{1}{4}}s^{-\frac{3}{4}}$
$\hat{a}$ $=ap_{\max}^{-\frac{1}{2}}s^{-\frac{1}{2}}$
$\hat{\jmath}$ $=jp_{\max}^{-\frac{3}{4}}s^{-\frac{1}{4}}$
@@
allows us to plot a three-dimensional fish diagram (see Figure 8). The contour surfaces of constant $\hat{T}$ values are indicated by the colored curves circulating around the fish. The conditions
@@
$\hat{\jmath}^{\star}=\sqrt[4]{\frac{1}{8}},\quad\hat{a}^{\star}=\sqrt{\frac{1}{8}},\quad\hat{v}^{\star}=\sqrt[4]{\frac{1}{32}},\quad\hat{T}^{\star}=\sqrt[4]{512}.$ (67)
@@
  • The bounds $\hat{v}_{\max}$, $\hat{a}_{\max}$ and $\hat{\jmath}_{\max}$ are now cutting planes.
  • If no bound is active (i.e. the solution of MTT is equivalent to the global minimum), $\dddot{x}(t)$ is “fully triangular”, and none of the orders has a cruising phase. Hence this kind of motion consists of 8 phases which is the minimum number of phases for a snap-controlled motion.
@@
  • This distinction can be made systematically by switching the conditions of (66) on or off independently from each other and counting how many combinations result (indeed, $2^{3}=8$).
@@
Figure 8: Three-dimensional plot of $\hat{T}$ as a function of $\hat{v}$, $\hat{a}$ and $\hat{\jmath}$ according to Equations (61, 65). The surfaces of constant $\hat{T}$ value are indicated by the colored curves (from orange for high $\hat{T}$ to dark purple for low $\hat{T}$), where the actual surface embraced by the curves is not shown for a better visibility of the entire figure. The three red curves frame the region of all admissible combinations of ($\hat{v},\hat{a},\hat{\jmath}$). They are obtained by combining two of the three conditions of (66), e.g., $\hat{T}_{1}=2\hat{T}_{2}$ and $\hat{T}_{2}=2\hat{T}_{3}$ fix the red solid curve. A single condition like $\hat{T}_{1}=2\hat{T}_{2}$ corresponds to a surrounding surface (not shown). The global minimum which is determined by the simultaneous consideration of all conditions of (66) is marked by the blue cross.
@@ 5.4 Systematics for Arbitrary Order
As already mentioned, such a motion has at least 8 phases and at most 15 phases. The click already mentioned at the end of Section 5.3 is the 5th order $x^{(5)}(t)$. Setting $c=x_{5}=\max\limits_{t\in[0,T]}|x^{(5)}(t)|$ as achieved click leads to
@@
Also even higher order, for example the yeet $x^{(6)}(t)$ and the pow $x^{(7)}(t)$, can be considered ([18] goes to 9th order.). We will continue with an arbitrary order $N$. The dimensional transformation to obtain $\hat{x}_{0}=\hat{x}_{N}=1$ is in general given by
$\hat{x}_{n}$ $=x_{n}x_{0}^{\frac{n}{N}-1}x_{N}^{-\frac{n}{N}}$ (73)
$\hat{T}$ $=Tx_{0}^{-\frac{1}{N}}x_{N}^{\frac{1}{N}}.$ (74)
@@
$\hat{x}^{\star}_{N-1}$ $=2^{\frac{1}{2N}(N-1)(2-N)}$ (75)
$\hat{x}^{\star}_{n}$ $=\hat{x}_{N-1}^{{\star}{N-n}}2^{\frac{1}{2}(n-N+1)(n-N+2)}$ (76)
@@
Clearly, $\hat{T}^{\star}_{n-1}=\hat{x}^{\star}_{N-1}$. Due to Equation (69) it is easy to find that
$\hat{T}^{\star}_{n}=2^{N-n-1}\hat{x}^{\star}_{N-1}$ (79)
@@
as a recursion for $h$ (with h(N1)=0h(N-1)=0). Then the solution is $h(n)=\frac{1}{2}(n-N+1)(n-N+2)$ which proves Equation (76). The condition $\hat{T}^{\star}_{0}=2\hat{T}^{\star}_{1}$ which is equivalent to $\frac{1}{\hat{x}^{\star}_{1}}=\hat{T}^{\star}_{1}$ finally determines $\hat{x}^{\star}_{N-1}$ by using Equations (76) and (78). Equation (78) is a direct implication. ∎
The Table 1 lists some examples.
N $\hat{T}^{\star}_{0}$ $\hat{v}^{\star}$ $\hat{a}^{\star}$ $\hat{\jmath}^{\star}$
@@ 4 $\sqrt[4]{512}$ $\sqrt[4]{\frac{1}{32}}$ $\sqrt{\frac{1}{8}}$ $\sqrt[4]{\frac{1}{8}}$
It can be seen from Equation (77) that the minimal time grows exponentially with the order of the motion [^9] While for the jerk-controlled motion $\hat{T}^{\star}_{0}\approx 3.17$, the case $N=6$ (yeet-controlled motion) and $N=7$ (pow-controlled motion) yield $\hat{T}^{\star}_{0}\approx 10.08$ and $\hat{T}^{\star}_{0}\approx 14.49$.. So the time intervals, where the highest order is non-vanishing, become rapidly shorter and the ratios $\frac{x_{n}}{x_{n+1}}$ for higher $n$ shrink.
[…]
$\hat{T}(\hat{z})=\frac{1}{\hat{z}_{1}}+\frac{\hat{z}_{1}}{\hat{z}_{2}}+\ldots+\frac{\hat{z}_{N-2}}{\hat{z}_{N-1}}+\hat{z}_{N-1}$ (84)
for $\hat{z}=(\hat{z}_{1},\ldots,\hat{z}_{N-1})\in\mathbb{R}^{N-1}$ (i.e. the time horizon) has no local minimum within the admissible region framed by the hypersurfaces encoded in $\hat{T}_{n-1}=2\hat{T}_{n}$. Thus, the solution of MTT is located at boundaries $\hat{w}_{n}$ or at surfaces $\hat{T}_{n-1}=2\hat{T}_{n}$. So $\hat{x}_{n}^{\star}=\hat{w}_{n}$ (active constraint) or there is a condition $\hat{T}_{n-1}=2\hat{T}_{n}$ for determining all $\hat{x}^{\star}_{n}$ which are not equal to a bound (i.e. $\hat{x}^{\star}_{n}<\hat{w}_{n}$). The formal algorithm to find the minimum-time trajectory reads as follows:
  1. 1.
    Transform the bounds $w_{n}$ to $\hat{w}_{n}$ according to Equation (73) within $x_{N}=w_{N}$ for $n=1,\ldots,N-1$. Set $\hat{w}_{0}=1$.
  2. 2.
    For $n=0,\ldots,N-1$ compute
    $\hat{T}_{n,\hat{w}}=\sum^{N-1}_{k=n}\frac{\hat{w}_{k}}{\hat{w}_{k+1}}.$ (85)
  3. 3.
    Collect all indices $n$ where $\hat{T}_{n-1,\hat{w}}<2\hat{T}_{n,\hat{w}}$, i.e. find the sets
    $B$ $=\left\{n\penalty 10000\ |\penalty 10000\ \hat{T}_{n-1,\hat{w}}<2\hat{T}_{n,\hat{w}}\wedge 0\leq n\leq N-1\right\}$ (86)
    $\bar{B}$ $\B\displaystyle=\left\{0,\ldots,N-1\right\}\backslash$ B (87)
  4. 4.
    For all $n\in\bar{B}$ set $\hat{x}^{\star}_{n}=\hat{w}_{n}$.
  5. 5.
    Solve the equation system
[…]
    for $\hat{z}\in\mathbb{R}^{n}$ with $n\in B$ and $\hat{T}_{n}(\hat{z})=\sum_{k=n}^{N-1}\frac{\hat{z}_{k}}{\hat{z}_{k-1}}$ where $\hat{z}_{0}=\hat{z}_{N}=1$ and $\hat{z}_{n}=\hat{x}^{\star}_{n}=\hat{w}_{n}$ for $n\in\bar{B}$.
This algorithm allows a direct computation of the reached values of velocity, acceleration, $\ldots$, without complicated case distinction. Since the $\hat{T}_{n}$ values follow directly the trajectory can be constructed directly as well.
@@ 6 Summary and Outlook
for jerk-controlled motions, is valid for any number of phases the motion can have. Furthermore, arbitrarily high orders can be considered. This allows for a reinterpretation and solution of well-known standard trajectory optimization problems (such as motions with minimum time or motions with minimum velocity). In particular, the condition $T_{n}=2T_{n+1}$ enables the determination of the active constraints for the respective case. It can also be estimated what effect considering higher orders has on the overall time horizon. Therefore, it seems surprising that – although the equation is not exceptionally complex and covers a large number of cases – it has not previously been presented in this form in the literature.
The equations for $T_{n}$ in combination with the conditions $T_{n}=2T_{n+1}$ opens up the possibility of direct calculation of the minimum-time motion.
Besides the practical case of application, it should be noted that motions with the minimum velocity are also energy-optimal in many cases, because a long phase without acceleration combined with a short phase of higher acceleration outweights motions with long phases of lower acceleration.
Clearly, many extensions exist. First, the zero boundary condition can be relaxed. Furthermore, it would be possible to consider other classes of trajectories (for example trajectories with circular blends of trigonometric trajectories). The expected results would then likely be considerably more unwieldy than those obtained here.
Another possibility – particularly mathematically interesting – lies in the limit $N\rightarrow\infty$ of infinite order. This raises questions of convergence (see the series in Equation (32) in Theorem 5) as well as questions concerning the relationship with $\mathcal{C}^{\infty}$ functions and mollifiers. Due to the infinitely large number of intervals in a finite time interval, it is also worthwhile to consider methods of fractal geometry (such as fractal strings, see [20]). However, all these investigations of interesting connections are left for future work.